State whether the following statements are true or false. Give reasons for your answers (1) The function f:R^3®R, given by f( x, y, z) =|x|+|y|+|z| is differentiable at (2, 3,-1).


1
Expert's answer
2020-03-20T12:14:25-0400

There exist neighbourhood U of (2, 3, -1) such that"\forall (x, y, z) \in U, f(x, y, z) = x + y - z" by definition of the absolute value.So, if "(x, y, z) \in U:\\" "\frac{\partial f}{\partial x} = \frac{\partial f}{\partial y} = 1\\
\frac{\partial f}{\partial z} = -1\\" Since function "g(x, y, z) = const" is continuous, all partial derivatives of f(x, y, z) are continuous at (2, 3, -1). Thus, f(x, y, z) is differentiable at (2, 3, -1).